OpenStudy (anonymous):

solve for s A=7s^(2)

OpenStudy (anonymous):

I get s=sqrtA/7

OpenStudy (anonymous):

or is it s=-sqrt A/7

OpenStudy (anonymous):

I don't think so. I think it goes like\[A=7s^2\rightarrow A/7=s^2\rightarrow \pm \sqrt{A/7}=s\]

OpenStudy (anonymous):

When you are taking the square root, you are leaving out the seven.

OpenStudy (anonymous):

A=7s^2

Directrix (directrix):

A=7s^(2) (1/7) A = (1/7) (7) (s^2) A/7 = s^2 ±√ (A/7) = √(s^2) ±√ (A/7) = s ±√(A) /√7 = s ±√(A) /√7 * 1 = s ±√(A) /√7 * √(7) / √(7) = s ±√(7 A) / [ √(7) * √(7 ] = s ±√(7 A) / 7 = s

OpenStudy (anonymous):

my choices are s=A/7 s=sqrt A/7 s=-sqrt A/7 or solutions are not given

Directrix (directrix):

 If A is a formula for area, then a negative value of s is not applicable. Of the options, that leaves s=sqrt A/7 from your list. My question is where are the grouping symbols? I've drawn the two quantities that sqrt A/7 could mean. The importance of grouping symbols in clarifying what is meant by an expressive cannot be overemphasized. So, post which option is on your paper or in your book and I will help you figure out what the answer is from among the options.